OpenStudy (anonymous):

john thunderhawk is making furniture he finds that to design a particular cabitnet and set up his shop to produce it cost $1400. then the labor and supplies to build each one are $400. so his cost,y, for x items is respentented by y=1400+400x. he can sell the cabinets for $600. so his revenue y is repesentive by y=600x. the solution tothis system of equations is his break even piont the point beyond which he starts making profit. How many cabinets must he sell to break even?

OpenStudy (anonymous):

i need help please

OpenStudy (anonymous):

he must sell 7 cabinets to break even

OpenStudy (anonymous):

1400+400=1800 -600=1200 cabinet 1 1200+400=1600 1600-600=1000 cabinet 2 1000+400=1400 1400-600=800 cabinet 3 800+400=1200 1200-600=600 cabinet 4 600+400=1000 1000-600=400 cabinet 5 400+400=800 800-600=200 cabinet 6 200+400=600 cabinet 7 600-600=0 cabinet 7

OpenStudy (anonymous):

thanks you for helping me

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

was my answer correct??

OpenStudy (anonymous):

yes